News

cecs_frontis
CECs, Centro de Estudios Científicos (Center for Scientific Studies) is a private, non-profit corporation, devoted to the development, promotion and diffusion of scientific research. CECs was founded in 1984 as the Center for Scientific Studies of Santiago and has since been directed by physicist Claudio Bunster.



Second DGA Expedition 2015 PDF Print E-mail

1.1

 

 

Monday 12th October

 

The wind is at last calm, but the “plateau” is now completely covered in cloud and it is snowing. All is well at the camp, according to radio reports from Rodrigo.

Reports from the group working on the beach of O'Higgins glacier reveal all to be well with them too. The weather here has been overcast, slightly windy, with no rainfall. During the morning, the group will try to take measurements in front of the tongue of O'Higgins glacier.

Meanwhile, Tobias and Marco remain at the Villa O'Higgins hostería, awaiting better weather conditions for radar measurements at the “plateau”.

Skies are overcast and there is almost no wind. These conditions create a significant loss of contrast, such that airborne radar measurements, it was decided, were best left for a better moment. During the afternoon today, Andrés and Ryan made water pressure measurements at O'Higgins Lake, facing O'Higgins glacier, while Jonathan, Sebastián and Jorge H. remained at the beach camp, to the south of the tongue of O'Higgins glacier. They are all well.

The group at the “nunatak” camp of O'Higgins glacier, managed to install three of the four component parts of the base of module 2. From the “plateau”, Rodrigo reported very cloudy skies, sporadic snow showers and light winds.

 

Sunday 11th October

 

Throughout the day today, we have accomplished several tasks. The Villa O'Higgins group (Tobias and Marco) made measurements with airborne LIDAR , thus completing the planned line measurements for the day at the “plateau”, covering a distance of 350 km.

The “nunatak” camp group, at O'Higgins glacier, completed installation of the base of module 1 and also settled the base of module 2 into its final position, pending attachment to the rock.

Team Jorge H, Sebastián and Ryan went to the tongue of the Chico glacier to install beacons, while Andrés and Jonathan made bathymetric measurements at O'Higgins Lake, facing the glacier.

 

   

 

  

 

Saturday 10th October

 

We had low temperatures during the morning, but skies were clear and there was no wind. Rodrigo informed us, at first light, of similar weather conditions at the “nunatak” camp of O'Higgins glacier. Based on this, we made an airborne LIDAR measurement (team: Tobias and Marco), between 09:45 and 13:00 hrs. During the latter part of the afternoon, from 13:45 to 16:00 hrs, another measurement was made. In total, the LIDAR team covered a vast ice distance of 650 km.

Andrés, Sebastián, Jorge H., Jonathan and Ryan prepared their cargos and transferred to the peninsula that separates the small arm of O'Higgins Lake. Jonathan and Ryan were flown to the peninsula, while Andrés, Sebastián and Jorge H. boarded the the CECs 2 vessel and crossed O'Higgins Lake, to the north side of the peninsula.

The group at the O'Higgins glacier camp began to build the living quarters. They made ready the bases of module 1 and had fixed a part of them to the “nunatak” by the end of the day.

Between 20:00 and 21:00 hrs, both the O'Higgins glacier group, as well as the peninsula group, reported that all was well.

 

Friday 9th October

 

A successful morning. Alejandro, Claudio and Jorge C. were transferred to the O'Higgins glacier camp and the living quarters modules were transported from the peninsula that separates the small arm fromf the main O'Higgins Lake. At the same time, a new Bubble Pulser test was performed at O'Higgins Lake.

In the afternoon, airborne LIDAR measurements were made over the Southern Patagonian Ice Field, covering a distance of 400 km. In addition, cargo was transported from Villa O'Higgins to the “plateau” camp. We also prepared supplies for the bathymetric measurement campaign.

We were in contact with Rodrigo throughout the day. Weather conditions have continued to be favourable, as of our last contact with him (approx. 20:30 hrs).

 

Thursday 8th October

 

Team Rodrigo, Carlos and Felipe remained at the high base camp at O'Higgins glacier, on the plateau of the Southern Patagonian Ice Field, resisting 20-30 knot winds. Weather conditions throughout the day have not permitted Claudio, Alejandro and Jorge C., at the plateau, to be transferred, to join the first group. Neither have we been able to transport cargo between the peninsula and the camp. Hopefully, weather conditions will improve tomorrow. Meanwhile, team Andrés, Jorge H, Jonathan, Sebastián and Ryan have been testing the Bubble Pulser at O'Higgins Lake, returning to Villa O'Higgins at around 17:00 hrs.

 

Wednesday 7th October

 

This morning, weather conditions are better than yesterday. At this precise moment, two research groups have left to work at the plateau and at Chico glacier, while the next two groups will head to the peninsula between Chico Lake and O'Higgins Lake and to the nunatak, where modules 1 and 2 will be installed, so we can make cargo flights between the two points. With respect to weather conditions, we hope for clearer weather tomorrow and that it lasts until Sunday.

 

    

 

Tuesday 6th October

 

In the morning, the radar antenna was taken, by CC-CSU helicopter, to the peninsula that separates the small arm of O'Higgins Lake. Marco and Jorge were on this flight, to respectively secure the antenna and to perform a geological survey. Meanwhile, Sebastián, Claudio and Alejandro were transferred via CC-AGG helicopter to the O'Higgins glacier to carry out work at the peripheral automatic weather station. Carlos and Jorge H. were also transferred to the plateau to work.

At approximately 15:30 hrs, Marco and Jorge returned safely to Villa O'Higgins, on board the CC-AGG. El helicopter landed to connect a load on the hook and then proceeded with its cargo flight to the peninsula. We hope that the group working in the periphery and the group on the plateau will both be evacuated during the afternoon (at approximately 17:00 hrs).

By the end of today we had successfully replaced the water pressure sensor at O'Higgins Lake. It had been destroyed by icebergs. The whole team is now back at Villa O'Higgins and enjoying a hearty dinner.

 

   

 

Monday 5th October

 

The weather this morning is similar to that we had yesterday, overcast but without rain. It was only possible to transport two loads, by helicopter, to the peninsula that separates Chico Lake from O'Higgins Lake. Part of the CECs team went to the runway to prepare the second transport cargo. Everyone else stayed at the base. The weather forecast for the next few days shows that tomorrow, Tuesday, the weather worsens, though there may be a window of better weather on Wednesday during which we might get to work at O'Higgins glacier.

The helicopters were back at 14 hrs, after having successfully delivered their cargo. Given the weather conditions, in particular the wind, flights were suspended for the rest of the day.

Part of the CECs team went together to the Villa O'Higgins landing strip to check camping equipment (2 Weatherhaven tents). The remainder of the team continued with research work at Villa O'Higgins base camp.

 

   

 

Sunday 4th October

 

 

Morning skies today had medium to high cloud cover and there was sporadic rainfall. Thus, all flight operations were suspended. Andrés, Jorge H., Jonathan and Ryan went to test the navigation system at O'Higgins Lake, at the Villa O'Higgins quay. The rest of the team stayed at base.

 

Friday 2nd October

 

We began to work at around 09:00 hrs, with the aim of performing cargo flights from Villa O’Higgins to the peninsula and surveying the “nunataks” to assess for installation of the CECs modules. Cloud cover was medium to high and there were high winds of more than 20 knots, making radar and “LIDAR” operations impossible.

Andrés, Sebastián, Jorge H and Jorge C. travelled in the CC-CSU helicopter while the CC-AGG helicopter took some of the module 1 equipment. At the same time the ground crew prepared two additional chinguillos con “gajos” (cargo slings made with “pieces”) from modules 1 and 2.

Both helicopters returned to Villa O’Higgins during the morning, to take a cargo sling with “pieces” and a sling/holder that had been put together the day before with bases and doors from module 2. In the field, Andrés and Sebastián performed a survey of the “nunataks” in the O’Higgins Corridor, while Jorge H. and Jorge C. carried out geological surveys and LIDAR land measurements.

 

  

 

Thursday 1st October

 

A day of opportunities: clear skies with only light winds. We decided to make airborne radar measurements from the CC-CSU helicopter, with Marco in charge of radar operations and Tobias navigating. The helicopter took off at approximately 10:00 hrs and returned to Villa O’Higgins at about 13:30 hrs. During this flight, many radar lines, with ice thickness, were measured.

On arrival at the base and considering the favourable weather conditions, from 15:00 hrs onwards we decided to proceed with LIDAR measurements for the rest of the day. Our helicopter returned at 17:30 hrs, having successfully carried out all programmed line measurements. At the same time, Jorge H, Ryan and Sebastián boarded the CC-AGG to perform work at the peripheral weather station of O’Higgins glacier and at Chico glacier.

 

Tuesday 29th September

 

Because of unfavourable weather conditions, today has come to a close without managing any flights. In general, today’s sky was completely overcast, with some light rain showers. In the intervals without rain, we installed the LIDAR. Poor weather conditions, snow blizzards and very poor visibility, were also reported from the plateau.

 

   

 

Monday 28th September

 

 

Weather conditions are good today, with clear skies and a few low scattered clouds (over the hills). The CC-AGG helicopter left at 11:30 am for work in the periphery of O’Higgins Glacier. This includes downloading data from the AWS, changing the automatic camera, checking the water level sensor and installing Ryan’s cameras. Jonathan, Jorge, Sebastian and Ryan were on board this flight.

The CC-CSU helicopter left at 12:00 pm, taking the radar antenna, mainly for East-West DGA line measurements. Navegator, Jorge Correa and radar operator, Marco, were on board this flight.

Claudio and Tobias reached Villa O'Higgins today, from Tortel, at around 3:00 pm. The CC-CSU helicopter, with radar antenna, also arrived at 3:00 pm, without difficulties. The East-West DGA lines were measured successfully, as well as measuring the tongue of O'Higgins Glacier. We made another radar flight (CC-CSU) that took off at 4:30 pm and landed at 7:00 pm, without problems.

The CC-AGG flight arrived back at 6:30 pm. All fine.

The group managed to complete the following work:

- Download O'Higgins AWS data.

- Change O'Higgins automatic camera.

- Install Ryan’s camera supports at Chico Glacier.

- Check the water level sensor.

 

    

 

Sunday 27th September

 

Our people are in Villa O’Higgins preparing cargos for the start of operations in this area of the Ice Field. At the plateau, Rodrigo and the others have suffered all day with the unpleasantness of white wind and the heavy accumulation of snow. During the night in Villa O’Higgins, the sky has cleared and the weather has improved, at least here - which has meant we have been able to enjoy the lunar eclipse to the full.

 

    

 

Friday 25th September

 

Throughout the morning today, the truck has been being loaded with all the storeroom goods, in order to transport everything to Villa O'Higgins and be there at 3 pm. The day dawned with some cloud cover, but skies are looking clear towards the south. However, Rodrigo has informed us that at this precise moment, weather conditions are bad at the plateau, with white wind and only 2m visibility. We thus plan to move only the antenna by helicopter to Villa O'Higgins.

All personnel have been moved to Villa O'Higgins without problems. The helo was able to reach Témpanos, where the AWS installation was completed and was left transmitting data with no problems. The modules were left operational and in very good shape. Claudio and Sebastián were then moved to Villa O’Higgins. On the way, Alejandro joined Rodrigo, Felipe and Carlos in Greve, to continue with work at that refuge.

 

Thursday 24th September

 

The helicopter flight that left at 12:40 pm was smooth and landed at 3:10 pm. People were moved without problems. Installation of the Témpano AWS has been possible, with functioning data transmission, and we await DGA confirmation. The modules have also been cleared.

There is complete cloud cover in Tortel. The pilot has informed us that though there has been a lot of cloud at the plateau, the contrast is poor for radar measurements, so today’s flights have been suspended. Geodesic GPS measurements of a point in Tortel were made during the afternoon. The storeroom has also been tidied, so that tomorrow we can all be moved, with cargo and antenna, to Villa O’Higgins.

 

    

 

Wednesday 23rd September

 

The LIDAR flight was planned for today, but was cancelled after detecting a small oil leak in the helicopter engine compartment. The second helicopter should arrive in the afternoon at about 4:00 pm. In the morning we communicated with the group on the plateau, who informed us that they have no plans to move anyone today.

The fine weather continues today in Tortel. It is a sunny day in the area, with a few scattered clouds over the hills. But at least there is no wind.

The second helicopter landed at 5:30 pm in Tortel, without difficulties. However, it’s already too late in the day for a radar flight. The weather conditions remain unchanged. The day has been partially cloudy and wind-free.

 

Tuesday 22nd September

 

The morning flight was uneventful. It left at 9:50 am and landed at 12:00 pm. Everyone stayed in Témpano. For now, the group will work until tomorrow, at 12:30 pm, when there is a scheduled group communication for coordination purposes. The group at Témpano is concentrating on some required repairs in the modules. After this, a radar flight was performed, leaving at 2:00 pm and landing at 4:30 pm. We did not manage to measure DGA lines because of local cloud cover, so instead we measured some more lines at Jorge Montt to increase the density of data in that area. Weather conditions have been generally good, with clear skies in the morning and some light winds, though some scattered clouds appeared over the hills during the afternoon.

 

    

 

Monday 21st September

 

At 8.40 am today, Carlos and Felipe left for the (Greve) plateau by helicopter. This flight landed at 10:45 am. Carlos and Felipe remained at Greve glacier, managing to get inside the CECs module, to find everything within in fine fettle. A second flight to Greve, carrying Rodrigo and Sebastián, left at 11:20 am. During the afternoon, we hope to make aerial radar measurements. Weather conditions remained good in the area, with completely clear skies throughout the day.

 

    

 

Sunday 20th September

 

We began the first flight radar measurements of the expedition. Marco and Tobías left on the 11:30 am flight. All instruments were functioning optimally and the weather continued to improve. Given the good weather conditions, we decided to launch a second radar flight that left at 3:50 pm and landed at 6:25 pm. A great day! A superb start to our radar campaign.

 

    

 

Friday 18th September

 

No flights have been made to the plateau or periphery today because of poor weather conditions. The Témpano (periphery) station indicated a relative humidity of almost 100% and it has been covered by low clouds over the hills. During the morning, test flights were made to carry out complete “checks” of the radar. All tests proved satisfactory. We can now rest and celebrate “The 18th” (Chilean Independence Day).

 

Thursday 17th September

 

Our flight left at 9:30 am, to “attack” the Témpano (periphery) with our team (Jorge, Sebastián and Alejandro). Today our mission is to download data and to perform maintenance of the automatic weather station (AWS), the automatic camera and the Témpano (periphery) pressure sensor.

The helicopter has landed at 3:25 pm. Tired but content, we successfully achieved all of today’s objectives despite cloudiness in the area throughout the day.

 

Tuesday 15th September

 

After days of irregular weather conditions, our helicopter finally took off at 8:40 am. Jorge and Sebastián will go to work on the camera at Jorge Montt while Felipe and Carlos will try to reach the Témpano (ice plateau) to check the weather station and download sensor data.

The helicopter landed at 13:00 pm. Our scientific missions at Jorge Montt were completed successfully, including changing the camera that contained photographic images up to August 15th. Because of cloud at the plateau, we did not manage to reach the Témpano.

There was complete cloud cover at the refuge, which meant we were unable to make a visual inspection of the module. Once again, we must postpone operations because of weather conditions.

 

    

 

Saturday 12th September

 

Finally! Weather conditions are looking favourable for work in the periphery. The Jorge Montt weather station indicates a relative humidity of 80%, a wind speed in the order of 2 to 4 m/s and no rainfall has so far been registered. The forecast indicates that the wind will increase as the day progresses. We decided to arrange for takeoff. The first flight left at 10:15 am. Our plan was for Jorge and Sebastián to reach the automatic camera at Jorge Montt glacier, while Alejandro and Claudio would work at the weather station and with the rain sensor in the same area.

…but all flights for today have been cancelled. During the first attempt to reach Jorge Montt (periphery), it was confirmed that the cloud cover was too low for helicopter operations, as well as discovering unexpected high winds at the higher altitude.

Today is not the day. We will wait…

 

Tuesday 8th September

 

The group of four scientists will remain at Villa O’Higgins until tomorrow. At the moment it is snowing in the area. We hope for some good weather over the next few days so that we can begin flight operations.

 

    

 

Monday 7th September

 

Today at 10:30 am, Alejandro, Marco, Sebastián and Carlos were transferred to Villa O'Higgins to unload the CECs module and the rest of the implements. Rain is predicted all day today. It is raining right now.

 

Sunday 6th September

 

The Second DGA Expedition 2015 began at 17:00 hrs, with our transport and load arriving in Bahía Tortel. All infrastructural elements and equipment arrived in good condition. Following the required standard checks, the parts of the buildable structure of the 20 MHz radar antenna were transferred from the airport warehouse.